# Learning DFAs from Positive Examples Only via Word Counting

### v1

@@ 5 Experiments
Our goal is now to use our new word-counting perspective to improve the state-of-the-art algorithm solving Task 2. All implementations can be found on GitHub.
Before we discuss our algorithms, let us describe the experimental setup that we have used.
@@
In order to compare our algorithms with the state-of-the-art symbolic algorithm developed in DBLP:conf/aaai/0002GBN0T23, we have used the inputs used in that paper to compare the performance of the symbolic algorithm and the counterexample guided algorithm developed in DBLP:conf/icgi/AvellanedaP18. These inputs are constituted of 28 samples of around 1000 words of lengths from 1 to 10. These words are obtained by sampling the languages of randomly generated small DFAs (of at most 10 states), with an alphabet of size 3. As in DBLP:conf/aaai/0002GBN0T23, we consider a timeout (TO) of 1000s to solve instances of Task 2 and Task 6. All experiments are done with a number of states $n$ between 1 and 8. For all samples, we have run the algorithms 10 times, and we use the mean runtime (with the standard deviation).
@@ 5.1 ILP algorithm solving Task 6
Let us describe on a high level how our ILP encoding works—the details can be found in Appendix C. Our encoding relies on two types of variables: binary variables, encoding the DFA and checking that it is in $\mathsf{Rec}(\mathcal{P},n)$; and integer variables which are used to count the number of accepted words. The inequalities involving the binary variables are actually very close to the logical constraints used in the SAT-encoding developed in DBLP:conf/aaai/0002GBN0T23. As for the integer variables, our encoding relies on Lemma 8: the inductive relation described in this lemma allows us to store in an integer variable the number of words of length at most 2n22n-2 accepted by the prospective DFA, which we can set as a minimization goal, subject to the other inequalities. Note that, to avoid having variable multiplication, we had to use the classical Big-M method griva2008linear.
We implemented our algorithm in Gurobi gurobi which we have used with a Python3 API. In Figure 2, we have plotted in blue dots the computation time of our ILP algorithm (y-axis) and the computation time of the symbolic algorithm (x-axis) on each 28 samples with $n=4$. Clearly, the ILP algorithm largely under-performs the symbolic algorithm. The situation is not significantly better for the case $n=6$ plotted as red stars [^2] We only plotted the samples for which the ILP algorithm did not reach a full timeout (i.e., all 10 attempts timeout) before $n=6$.. Overall, our ILP algorithm does not stand the comparison with the symbolic algorithm. We discuss below three possible reasons for this significant runtime difference.
@@
First, from a theoretical standpoint, the symbolic algorithm could asymptotically take exponentially many steps to terminate. This is why the ILP algorithm has better theoretical guarantees. However, in practice, the symbolic algorithm does not take many steps to terminate (around 20, in the worst case), and thus greatly outperforms our ILP algorithm. There could be some kinds of samples for which the symbolic algorithm takes much more steps to conclude, in which case the runtime difference could be much smaller.
@@ 5.2 A pre-processing algorithm
The goal of this heuristic algorithm is to output a DFA $\mathcal{A}\in\mathsf{Rec}(\mathcal{P},n)$ that accepts as few words of length at most 2n22n-2 as possible. The general idea of the algorithm is to randomly search among DFAs, and pick one that minimizes the number of accepted words. However, it is not clear how to randomly search among DFAs in $\mathsf{Rec}(\mathcal{P},n)$. Therefore, our heuristic algorithm focuses on transition systems instead. A transition system can be seen as a DFA without starting or final states, i.e., it is described by a tuple $T=(Q,\Sigma,\delta)$ with $\delta\colon Q\times\Sigma\to Q$. Consider then a transition system $T=(Q,\Sigma,\delta)$ with $|Q|\leq n$ and a starting state $q\in Q$. It is clear that choosing $F_{T,q}:=\{\delta^{*}(q,w)\mid w\in\mathcal{P}\}$ minimizes the number of words accepted by the DFA $\mathcal{A}_{T,q}:=(Q,\Sigma,\delta,q,F_{T,q})$ while ensuring that $\mathcal{A}_{T,q}\in\mathsf{Rec}(\mathcal{P},n)$. Following, given a transition system $T$, we define the score of the transition system $T$ as follows: $\min_{q\in Q}|\mathcal{L}(\mathcal{A}_{T,q})\cap\Sigma^{\leq 2n-2}|$. We naturally extends the notion of score to DFAs. Our heuristic algorithm then proceeds as follows:
@@
    From the current transition system, look through all possible change of a single transition. Make the change that improves the score the most, if one exists. Otherwise, stop. (Theoretically, this could take exponentially many steps to terminate, in practice, around 30 steps suffice.)
@@
This heuristic algorithm has two hyper-paraneters: $\mathsf{InitRand}$ and $\mathsf{NbRun}$. From a few experiments conducted on sample 02 (with $\mathsf{InitRand}\in\{10,100,1000\}$ and $\mathsf{NbRun}\in\{10,20,50,100\}$), we have chosen somewhat arbitrarily $\mathsf{InitRand}=100$ and $\mathsf{NbRun}=50$ as a compromise between running time and efficiency, though it is very likely that better values could be chosen (even for sample 02).
@@
There are two main positive takeaways from these results. First, there are three samples (02, 05, 08) where our algorithm (heuristic+symbolic*) significantly outperforms the symbolic algorithm. This is a promising result with a portfolio approach in mind where both the heuristic+symbolic* and symbolic algorithms would be run in parallel. Second, there are five samples (02, 05, 06, 08, 17) where the computation time of the symbolic* algorithm is much lower than that of the symbolic algorithm. This result is promising for harder instances where the runtime of the heuristic algorithm becomes more-or-less negligible compared to the runtime of the symbolic algorithm.
On the negative side, on the five other samples, the symbolic* algorithm does not outperform the symbolic algorithm, and even under-performs it for sample 27. This shows that a lower starting score (i.e., score of a starting DFA) does not always mean a better performance of the symbolic algorithm. To better identify how the starting score impacts the runtime of the symbolic algorithm, we have represented relevant data in Figures 4 and 5.
In Figure 4, each data point corresponds to a run of the symbolic* algorithm [^3] To compare the scores, we have only depicted the samples where the computation of the heuristic+symbolic* algorithm is done with $n=8$, thus we excluded sample 08.: in the x-axis, the starting score is depicted, in the y-axis is depicted the quotient of the runtime of the symbolic* algorithm and the mean of the runtime of the symbolic algorithm on that sample. In Figure 5, we have depicted the Pearson corelation coefficient between the runtime and the starting score. For each sample, we have included the score and runtime data of both the symbolic [^4] On all 10 runs, the starting score is $|\{a,b,c\}^{\leq 14}|$. and the symbolic* algorithms; while Total takes into account all sample results (except sample 08) at once. For Total*, we have computed the corelation between the quotient of the runtime of the symbolic(*) algorithms and the starting score, which corresponds to the data depicted in Figure 4.
From these figures, we can extract the following facts: a lower score is likely correlated to a lower runtime, and a very low score probably means a (very) short runtime. These facts tend to indicate that improving the capacity of our heuristic algorithm at finding DFAs with (very) low score (when possible) may have the potential to greatly lower the runtime of the heuristic+symbolic* algorithm.
Sample 02 05 06 08 15 17
r 0.98 0.94 0.78 0.74 -0.42 0.66
Sample 21 24 27 28 Total Total*
r 0.41 -0,12 -0.33 0.06 0.6 0.78
Figure 5: The Pearson correlation coefficient between the score and the runtime.
@@ 6 Conclusion
In this paper, we have studied the task of learning DFAs from positive examples only with a novel word-counting perspective. We have been able to show interesting theoretical results, in particular the first complexity result on a decision problem directly related to this task. On the experimental side, our ILP algorithm performs poorly, but our heuristic algorithm shows promising results as a pre-processing tool to use prior to running the symbolic algorithm. Notably, this algorithm shows that transition systems are particularly well-suited for this learning from positive examples task.
Despite our few experiments, how the starting score impacts the runtime of the symbolic algorithm remains rather cryptic. We believe that a better understanding of this impact could be valuable in the design of a heuristic algorithm more efficient at finding good starting DFAs.

### v2

@@ 5 Experiments
Our goal is now to use our new word-counting perspective to improve the state-of-the-art algorithm solving Task 2. All of our code and data is available on GitHub (cf. the top page). Before we discuss our algorithms, let us describe the experimental setup that we have used.
@@
In order to compare our algorithms with the state-of-the-art symbolic algorithm developed by Roy et al. (2023), we have used the inputs used in that paper to compare the performance of the symbolic algorithm and the counterexample guided algorithm developed by Avellaneda and Petrenko (2018). These inputs are constituted of 28 samples of around 1000 words of lengths from 1 to 10. These words are obtained by sampling the languages of randomly generated small DFAs (of at most 10 states), with an alphabet of size 3. As in [DBLP:conf/aaai/0002GBN0T23], we consider a timeout (TO) of 1000s to solve instances of Task 2 and Task 6. All experiments are done with a number of states $n$ between 1 and 8. For all samples, we have run the algorithms 10 times, and we use the mean runtime (with the standard deviation).
@@ 5.1 ILP algorithm solving Task 6
Let us describe on a high level how our ILP encoding works—the details can be found in the extended version [bordais2025learningdfaspositiveexamples]. Our encoding relies on two types of variables: binary variables, encoding the DFA and checking that it is in $\mathsf{Rec}(\mathcal{P},n)$; and integer variables which are used to count the number of accepted words. The inequalities involving the binary variables are actually very close to the logical constraints used in the SAT-encoding developed by [DBLP:conf/aaai/0002GBN0T23] (2023). As for the integer variables, our encoding relies on Lemma 8: the inductive relation described in this lemma allows us to store in an integer variable the number of words of length at most 2n22n-2 accepted by the prospective DFA, which we can set as a minimization goal, subject to the other inequalities. Note that, to avoid having variable multiplication, we had to use the classical Big-M method [griva2008linear].
We implemented our algorithm in Gurobi [gurobi] which we have used with a Python3 API. In Figure 2, we have plotted in blue dots the computation time of our ILP algorithm (y-axis) and the computation time of the symbolic algorithm (x-axis) on each 28 samples with $n=4$. Clearly, the ILP algorithm largely under-performs the symbolic algorithm. The situation is not significantly better for the case $n=6$ plotted as red stars [^2] We only plotted the samples for which the ILP algorithm did not reach a full timeout (i.e., all 10 attempts timeout) before $n=6$.. Overall, our ILP algorithm does not stand the comparison with the symbolic algorithm. We discuss below three possible reasons for this significant runtime difference.
@@
First, from a theoretical standpoint, the symbolic algorithm could asymptotically take exponentially many steps to terminate. However, in practice, the symbolic algorithm does not take many steps to terminate (around 20, in the worst case), and thus greatly outperforms our ILP algorithm. There could be some kinds of samples for which the symbolic algorithm takes much more steps to conclude, in which case the runtime difference could be much smaller.
@@ 5.2 A pre-processing algorithm
The goal of this heuristic algorithm is to output a DFA $\mathcal{A}\in\mathsf{Rec}(\mathcal{P},n)$ that accepts as few words of length at most 2n22n-2 as possible. The general idea of the algorithm is to randomly search among DFAs, and pick one that minimizes the number of accepted words. However, it is not clear how to randomly search among DFAs in $\mathsf{Rec}(\mathcal{P},n)$. Therefore, our heuristic algorithm focuses on transition systems instead. A transition system can be seen as a DFA without starting or final states, i.e., it is described by a tuple $T=(Q,\Sigma,\delta)$ with $\delta\colon Q\times\Sigma\to Q$. Consider then a transition system $T=(Q,\Sigma,\delta)$ with $|Q|\leq n$ and a starting state $q\in Q$. It is clear that choosing $F_{T,q}:=\{\delta^{*}(q,w)\mid w\in\mathcal{P}\}$ minimizes the number of words accepted by the DFA $\mathcal{A}_{T,q}:=(Q,\Sigma,\delta,q,F_{T,q})$ while ensuring that $\mathcal{A}_{T,q}\in\mathsf{Rec}(\mathcal{P},n)$. Following, given a transition system $T$, we define the score of the transition system $T$ as follows: $\min_{q\in Q}|\mathcal{L}(\mathcal{A}_{T,q})\cap\Sigma^{\leq 2n-2}|$. We naturally extend the notion of score to DFAs. Our heuristic algorithm then proceeds as follows:
@@
    From the current transition system, look through all possible changes of a single transition. Make the change that improves the score the most, if one exists. Otherwise, stop. (Theoretically, this could take exponentially many steps to terminate, in practice, around 30 steps suffice.)
@@
This heuristic algorithm has two hyper-parameters: $\mathsf{InitRand}$ and $\mathsf{NbRun}$. From several experiments conducted on sample 02 (with $\mathsf{InitRand}\in\{10,100,1000\}$ and $\mathsf{NbRun}\in\{10,20,50,100\}$), we have chosen $\mathsf{InitRand}=100$ and $\mathsf{NbRun}=50$ as a compromise between running time and efficiency, though it is plausible that better values could be chosen (even for sample 02).
@@
There are two main positive takeaways from these results. First, there are three samples (02, 05, 08) where our algorithm (heuristic+symbolic*) significantly outperforms the symbolic algorithm. This is a promising result with a portfolio approach in mind where both the heuristic+symbolic* and symbolic algorithms would run in parallel. Second, there are five samples (02, 05, 06, 08, 17) where the computation time of the symbolic* algorithm is much lower than that of the symbolic algorithm. This result is promising for harder instances where the runtime of the heuristic algorithm becomes more-or-less negligible compared to the runtime of the symbolic algorithm.
On the negative side, on the five other samples, the symbolic* algorithm does not outperform the symbolic algorithm, and even under-performs it for sample 27. This shows that a lower starting score (i.e., score of a starting DFA) does not always mean a better performance of the symbolic algorithm. To better identify how the starting score impacts the runtime of the symbolic algorithm, we have represented relevant data in Figure 4.
In this Figure, each data point corresponds to a run of the symbolic* algorithm [^3] To compare the scores, we have only depicted the samples where the computation of the heuristic+symbolic* algorithm is done with $n=8$, thus we excluded sample 08.: in the x-axis, the starting score is depicted, in the y-axis is depicted the quotient of the runtime of the symbolic* algorithm and the mean of the runtime of the symbolic algorithm on that sample.
From this figure, we can extract the following facts: when starting from a very low score, the runtime ratio is small (as exemplified for samples 02, 05, 17 and also possibly 27); and when starting from higher scores, the impact of the score on the runtime is hard to see. These facts may suggest that the significant differences on the various samples may mostly come from the fact that for some samples the heuristic algorithm has found a starting DFA with a (very) low score, while it is not the case for the other samples. Following, improving the capacity of our heuristic algorithm at finding DFAs with (very) low score (when possible) may have the potential to greatly lower the runtime of the heuristic+symbolic* algorithm. Nonetheless, how the starting score influences the runtime of the SAT solver remains rather cryptic despite our experiments, and further investigating this relationship seems warranted to improve the efficiency of the minimizing-score heuristic approach.
@@ 6 Conclusion
In this paper, we have studied the task of learning DFAs from positive examples only with a novel word-counting perspective. The core of our studies is theoretical: we have shown the first complexity result on a decision problem directly related to this task, and we have exhibited a new algorithm solving this task with better theoretical guarantees than the state-of-the-art. On the experimental side, our ILP algorithm performs poorly, but our heuristic algorithm shows promising results as a pre-processing tool to use prior to running the symbolic algorithm. Notably, this algorithm shows that transition systems are particularly well-suited for handling this learning from positive examples task.
Our main takeaway is that a better understanding (both theoretically and experimentally) of the relationship between the language minimality of DFAs and their number of accepted words would constitute a valuable insight in the design of efficient algorithms learning DFAs from positive examples only.
Acknowledgements. This work has been financially supported by Deutsche Forschungsgemeinschaft, DFG Project number 434592664.